OpenStudy (anonymous):

When a product is raised to a power, it is equal to the product of each factor raised to that power. Show that this rule does not apply for a sum or difference to a power using the problem (5 – 3)2. You may indicate an exponent in your answer with ^. For example, 3x2 as 3x^2.

OpenStudy (anonymous):

(5-3)^2 = (8)^2 = 64 5^2 - 3^2 = 25 - 9 = 16 obviously 64 does not equal 16 so distribution of the exponents when there is addition/subtraction is not the same

OpenStudy (anonymous):

sweet thanks